# Bidiagonal matrix factorisations related to multiple orthogonal polynomials

### v1

@@ 2. $\bullet$
Therefore, using (4.21), we find that
$g_{1}(t)-g_{0}(t)=\sum_{n=0}^{\infty}q_{k}\left[x^{(r+1)(n+1)}\right]\,t^{n+1}=\sum_{n=0}^{\infty}\alpha_{0}\,q_{r}\left[x^{(r+1)n+r}\right]\,t^{n+1}=\alpha_{0}\,t\,g_{r+1}(t),$ (4.23)
@@
$g_{k+1}(t)-g_{k}(t)$ $=\sum_{n=1}^{\infty}\left(q_{k}\left[x^{(r+1)n+k+1}\right]-q_{k-1}\left[x^{(r+1)n+k}\right]\right)\,t^{n}$ (4.24)
$=\alpha_{k}\sum_{n=0}^{\infty}q_{k+r}\left[x^{(r+1)(n+1)+k}\right]\,t^{n+1}=\alpha_{k}\,t\,g_{k+r+1}(t)\quad\text{for all }k\geq 1.$
Hence, $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ defined by (4.18) or, equivalently, by (4.21) satisfies (4.3). ∎
@@ 5. Case study: Jacobi-Piñeiro polynomials
In this section, we present a bidiagonal factorisation for the Hessenberg recurrence matrix of the Jacobi-Piñeiro multiple orthogonal polynomials on the step-line, obtained using both methods developed in Sections 3 and 4. At the end of this section, we also give a bidiagonal matrix factorisation associated with the multiple Laguerre polynomials of first kind as a limiting case of the Jacobi-Piñeiro polynomials.
For a positive integer $r$ and parameters $a_{1},\cdots,a_{r},b>-1$ such that $a_{i}-a_{j}\not\in\mathbb{Z}$ whenever $i\neq j$, we consider the measures $\left(\mu_{1},\cdots,\mu_{r}\right)$ supported on the interval $(0,1)$ with densities $\mathrm{d}\mu_{j}(x)=x^{a_{j}}(1-x)^{b}\,\mathrm{d}x$, whose moments are
$\int_{0}^{1}x^{n}\mathrm{d}\mu_{j}(x)=\int_{0}^{1}x^{a_{j}+n}(1-x)^{b}\,\mathrm{d}x=\mathrm{B}(a_{j}+n+1,b+1)\quad\text{for all }n\in\mathbb{N}\text{ and }1\leq j\leq r,$ (5.1)
@@
and $\Gamma(z)$ is Euler’s Gamma function.
The multiple orthogonal polynomials with respect to $\left(\mu_{1},\cdots,\mu_{r}\right)$ are the well-known Jacobi-Piñeiro polynomials introduced in [21]. Explicit expressions for the type I and type II Jacobi-Piñeiro polynomials with respect to an arbitrary number of weights can be found in [7] and [4], respectively.
@@ 5.1. Bidiagonal factorisation linked to branched continued fractions
Here, we apply our method to construct bidiagonal matrix factorisations associated with multiple orthogonal polynomials via their connection with branched continued fractions to the Jacobi-Piñeiro polynomials. Firstly, we need to find the sequence $g_{k}(t)$ satisfying (4.2) and (4.3), where $v_{1},\cdots,v_{r}$ are the functionals with the moments in (5.1) normalised so that $v_{j}\left[1\right]=1$ for all $1\leq j\leq r$. Then,
$v_{j}\left[x^{n}\right]=\int_{0}^{1}\frac{x^{a_{j}+n}(1-x)^{b}\,\mathrm{d}x}{\mathrm{B}\left(a_{j}+1,b+1\right)}=\frac{\mathrm{B}\left(a_{j}+n+1,b+1\right)}{\mathrm{B}\left(a_{j}+1,b+1\right)}=\frac{\left(a_{j}+1\right)_{n}}{\left(a_{j}+b+2\right)_{n}}\quad\text{for all }n\in\mathbb{N}\text{ and }1\leq j\leq r,$ (5.3)
where $\left(z\right)_{n}$ is the Pochhammer symbol defined by $\left(z\right)_{0}=1$ and $\left(z\right)_{n}=z(z+1)\cdots(z+n-1)$ for $n\geq 1$. Note that, defining $v_{1},\cdots,v_{r}$ by its moments in (5.3), we can replace the condition $a_{1},\cdots,a_{r},b>-1$ by the weaker condition that $a_{1},\cdots,a_{r},a_{1}+b+1,\cdots,a_{r}+b+1$ cannot be negative integers.
We define the sequence $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ by
$g_{k}(t)=(1-t)^{b}{}_{r+1}F_{r}\left(\left.\begin{matrix}a^{\,\prime}_{k}+b+1,\cdots,a^{\,\prime}_{k+r}+b+1\vspace*{2 pt}\\ a^{\,*}_{k+1}+b+1,\cdots,a^{\,*}_{k+r}+b+1\end{matrix}\,\right|\,t\right),$ (5.4)
where
  1. $\bullet$
    $a^{\,\prime}_{(r+1)n+j}=a_{j}+n$ for any $n\in\mathbb{N}$ and $0\leq j\leq r$, with $a_{0}=-1$, and
  2. $\bullet$
    $a^{\,*}_{rn+j}=a_{j}+n$ for any $n\in\mathbb{N}$ and $1\leq j\leq r$.
The initial elements of $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ are
$g_{0}(t)=(1-t)^{b}{}_{1}F_{0}\left(\left.\begin{matrix}b\\ -\end{matrix}\,\right|\,t\right)=1,$ (5.5)
and
$g_{k}(t)=(1-t)^{b}{}_{2}F_{1}\left(\left.\begin{matrix}a_{k}+b+1,b+1\\ a_{k}+b+2\end{matrix}\,\right|\,t\right)={}_{2}F_{1}\left(\left.\begin{matrix}a_{k}+1,1\\ a_{k}+b+2\end{matrix}\,\right|\,t\right)=\sum_{n=0}^{\infty}\frac{\left(a_{k}+1\right)_{n}}{\left(a_{k}+b+2\right)_{n}}\,t^{n}$ (5.6)
for $\text{for }1\leq k\leq r.$
To obtain our final expressions in (5.6), we use [13, Eq. 15.8.1]. Furthermore, the hypergeometric series ${}_{p}F_{q}$ satisfies the three-term contiguous relation (see [20, Lemma 14.1]):
${}_{p}F_{q}\left(\left.\begin{matrix}\vec{c}+\vec{e}_{i}\\ \vec{d}+\vec{e}_{j}\end{matrix}\,\right|\,t\right)-{}_{p}F_{q}\left(\left.\begin{matrix}\vec{c}\\ \vec{d}\end{matrix}\,\right|\,t\right)=\,\dfrac{\left(d_{j}-c_{i}\right)\prod\limits_{k=1,\,k\neq i}^{p}c_{k}}{\left(d_{j}+1\right)\prod\limits_{l=1}^{q}d_{l}}\;t\,{}_{p}F_{q}\left(\left.\begin{matrix}\vec{c}+\vec{1}-\vec{e}_{i}\\ \vec{d}+\vec{1}\end{matrix}\,\right|\,t\right),$ (5.7)
where $\vec{c}=\left(c_{1},\cdots,c_{p}\right)\in\mathbb{N}^{p}$, $\vec{d}=\left(d_{1},\cdots,d_{q}\right)\in\mathbb{N}^{q}$, $\vec{e}_{i}$ is the vector whose entries are all 00 except the $i^{\textrm{th}}$-entry which is equal to $1$, and $\vec{1}$ is the vector whose entries are all equal to $1$.
As a consequence, the sequence $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ in (5.4) satisfies the recurrence relation
@@ 2. $\bullet$
$\alpha_{n}=\frac{\left(a^{\,*}_{n+1}-a^{\,\prime}_{n}\right)\prod\limits_{j=1}^{r}\left(a^{\,\prime}_{n+j}+b+1\right)}{\prod\limits_{j=1}^{r+1}\left(a^{\,*}_{n+j}+b+1\right)}\quad\text{for any }n\in\mathbb{N},$ (5.9)
where $a^{\,\prime}_{n}$ and $a^{\,*}_{n}$ are the same as in (5.4). Writing $n=(r+1)(rm+k)+i$, with $m\in\mathbb{N}$, $10\leq k\leq r-$ 1 and $0\leq i\leq r$, we can write the coefficients $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ more explicitly. Then, using Theorem 4.1, we obtain the following result.
Theorem 5.1.
@@
The Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials on the step-line, multiple orthogonal with respect to $(\mu_{1},\cdots,\mu_{r})$ given by (5.1), admits the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$, as a product of bidiagonal matrices of the form in (1.2) with nontrivial entries $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ such that, for all $m\in\mathbb{N}$, $10\leq k\leq r-$ 1 and $0\leq i\leq r$, and setting $a_{0}=-1$, we have
$\alpha_{(r+1)(rm+k)+i}=\frac{\displaystyle\left(a_{k+i+1}-a_{i}+m\right)\prod_{j=0}^{i-1}\left(a_{j}+b+rm+k+2\right)\prod_{j=i+1}^{r}\left(a_{j}+b+rm+k+1\right)}{\displaystyle\prod_{j=1}^{k+i+1}\Big{(}a_{j}+b+(r+1)m+k+2\Big{)}\prod_{j=k+i+1}^{r}\Big{(}a_{j}+b+(r+1)m+k+1\Big{)}}\text{ if }k+i\leq r-1,$ (5.10)

### v2

@@ 4. Bidiagonal factorisations linked to branched continued fractions
As a result, we find that
$g_{1}(t)-g_{0}(t)=\sum_{n=0}^{\infty}q_{0}\left[x^{(r+1)(n+1)}\right]\,t^{n+1}=\sum_{n=0}^{\infty}\alpha_{0}\,q_{r}\left[x^{(r+1)n+r}\right]\,t^{n+1}=\alpha_{0}\,t\,g_{r+1}(t),$ (4.20)
@@
$g_{k+1}(t)-g_{k}(t)$ $=\sum_{n=0}^{\infty}\bigg(q_{k}\left[x^{(r+1)(n+1)+k+1}\right]-q_{k-1}\left[x^{(r+1)(n+1)+k}\right]\bigg)\,t^{n+1}$ (4.21)
$=\sum_{n=0}^{\infty}\alpha_{k}\,q_{k+r}\left[x^{(r+1)(n+1)+k}\right]\,t^{n+1}=\alpha_{k}\,t\,g_{k+r+1}(t)\quad\text{for all }k\geq 1.$
Hence, $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ defined by (4.14) or, equivalently, by (4.18) satisfies ((c)). ∎
@@ 5. Case study: Jacobi-Piñeiro polynomials
In this section, we present a bidiagonal factorisation of the Hessenberg recurrence matrix of the Jacobi-Piñeiro multiple orthogonal polynomials on the step-line, obtained using both methods developed in Sections 3 and 4. At the end of the section, we give a bidiagonal matrix factorisation associated with the multiple Laguerre polynomials of first kind as a limiting case of the Jacobi-Piñeiro polynomials.
For a positive integer $r$ and parameters b>1b>-1 and $a_{1},\cdots,a_{r}>0$ such that $i\neq j$ implies $a_{i}-a_{j}\not\in\mathbb{Z}$, we consider the measures $\left(\mu_{1},\cdots,\mu_{r}\right)$ supported on the interval $(0,1)$ with densities $\mathrm{d}\mu_{j}(x)=x^{a_{j}-1}(1-x)^{b}\,\mathrm{d}x$, whose moments are
$\int_{0}^{1}x^{n}\mathrm{d}\mu_{j}(x)=\int_{0}^{1}x^{a_{j}+n-1}(1-x)^{b}\,\mathrm{d}x=\mathrm{B}(a_{j}+n,b+1)\quad\text{for all }n\in\mathbb{N}\text{ and }1\leq j\leq r,$ (5.1)
@@
and $\Gamma(z)$ is Euler’s Gamma function. The multiple orthogonal polynomials with respect to $\left(\mu_{1},\cdots,\mu_{r}\right)$ are the well-known Jacobi-Piñeiro polynomials introduced in [23].
Explicit expressions for the type I and type II Jacobi-Piñeiro polynomials with respect to an arbitrary number of weights can be found in [6] and [5], respectively. These expressions involve hypergeometric series. For $p,q\in\mathbb{N}$, the hypergeometric series $_{p}F_{q}$ (see [3, 19, 13]) is defined by
$_{p}F_{q}\left(\left.\begin{matrix}a_{1},\cdots,a_{p}\\ b_{1},\cdots,b_{q}\end{matrix}\,\right|\,z\right)=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\cdots\left(a_{p}\right)_{n}}{\left(b_{1}\right)_{n}\cdots\left(b_{q}\right)_{n}}\,\frac{z^{n}}{n!}.$ (5.3)
where $\left(z\right)_{n}$ is the Pochhammer symbol defined by $\left(z\right)_{0}=1$ and $\left(z\right)_{n}=z(z+1)\cdots(z+n-1)$ for $n\geq 1$.
When $p\geq 1$, $a_{1}=-n$ for some $n\in\mathbb{N}$, $a_{2},\cdots,a_{p}\not\in\{0,-1,\cdots,1-n\}$ and $b_{1},\cdots,b_{q}\not\in\{0,-1,\cdots,-n\}$, (5.3) is a terminating series and defines a polynomial of degree $n$. Otherwise, we require that $b_{1},\cdots,b_{q}\not\in$-$\mathbb{N}$ and we treat (5.3) as a formal power series, without concerning about its convergence.
@@ 5.1. Bidiagonal factorisation linked to branched continued fractions
To construct a bidiagonal matrix factorisation of the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials on the step-line via their connection with branched continued fractions, we need to find the sequence $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfying ((c))-((c)), where $v_{1},\cdots,v_{r}$ are the functionals with the moments in (5.1) normalised so that $v_{j}\left[1\right]=1$ for all $1\leq j\leq r$, i.e.,
$v_{j}\left[x^{n}\right]=\int_{0}^{1}\frac{x^{a_{j}+n-1}(1-x)^{b}\,\mathrm{d}x}{\mathrm{B}\left(a_{j}+1,b+1\right)}=\frac{\mathrm{B}\left(a_{j}+n,b+1\right)}{\mathrm{B}\left(a_{j},b+1\right)}=\frac{\left(a_{j}\right)_{n}}{\left(a_{j}+b+1\right)_{n}}\quad\text{for all }n\in\mathbb{N}\text{ and }1\leq j\leq r.$ (5.4)
Note that, defining $v_{1},\cdots,v_{r}$ by its moments in (5.4), we can replace the conditions $a_{1},\cdots,a_{r}>0$ and b>1b>-1 by the weaker condition that $a_{1},\cdots,a_{r},a_{1}+b+1,\cdots,a_{r}+b+1$ cannot be non-positive integers.
Theorem 5.1.
For $a_{1},\cdots,a_{r}\in\mathbb{C}$ such that $a_{1},\cdots,a_{r},a_{1}+b+1,\cdots,a_{r}+b+1$ cannot be non-positive integers, let $\left(g_{n}(t)\right)_{n\in\mathbb{N}}$ be the sequence defined by
$g_{n}(t)=(1-t)^{b}\,{}_{r+1}F_{r}\left(\left.\begin{matrix}b+n_{0}^{\,\prime},a_{1}+b+n_{1}^{\,\prime},\cdots,a_{r}+b+n_{r}^{\,\prime}\vskip 2.0pt\\ a_{1}+b+n_{1},\cdots,a_{r}+b+n_{r}\end{matrix}\,\right|\,t\right)\quad\text{for all }n\in\mathbb{N},$ (5.5)
where $\vec{n}=\left(n_{1},\cdots,n_{r}\right)$ and $\vec{n}^{\,\prime}=\left(n_{0}^{\,\prime},\cdots,n_{r}^{\,\prime}\right)$ are the elements of the step-line of $\mathbb{N}^{r}$ and $\mathbb{N}^{r+1}$, respectively, such that $\left|\vec{n}\right|=\left|\vec{n}^{\,\prime}\right|=n$, i.e., $n_{j}=\left\lceil\frac{{n-j+1}}{r}\right\rceil$ for all $1\leq j\leq r$ and $n_{i}^{\,\prime}=\left\lceil\frac{{n-i}}{r+1}\right\rceil$ for all $0\leq i\leq r$. Then:
  • the initial elements of $\left(g_{n}(t)\right)_{n\in\mathbb{N}}$ are $g_{0}(t)=1$ and
    $g_{n}(t)=\sum_{k=0}^{\infty}\frac{\left(a_{n}\right)_{k}}{\left(a_{n}+b+1\right)_{k}}\,t^{k}\quad\text{for }1\leq n\leq r.$ (5.6)
  • $\left(g_{n}(t)\right)_{n\in\mathbb{N}}$ satisfies the recurrence relation
@@
    $\alpha_{n}=\frac{\left(a_{j(n)}-a_{i(n)}+\left\lfloor\dfrac{n}{r}\right\rfloor-\left\lfloor\dfrac{n}{r+1}\right\rfloor\right)\prod\limits_{i=0}^{i(n)-1}\left(a_{i}+b+\left\lfloor\dfrac{n}{r+1}\right\rfloor+1\right)\prod\limits_{i=i(n)+1}^{r}\left(a_{i}+b+\left\lfloor\dfrac{n}{r+1}\right\rfloor\right)}{\prod\limits_{j=1}^{j(n)}\left(a_{j}+b+\left\lfloor\dfrac{n}{r}\right\rfloor+1\right)\prod\limits_{j=j(n)}^{r}\left(a_{j}+b+\left\lfloor\dfrac{n}{r}\right\rfloor\right)},$ (5.8)
    for all $n\in\mathbb{N}$, with $a_{0}=0$, $i(n)=[n]_{r+1}$, and $j(n)=[n]_{r}+1$, where $n]_{k}$ denotes the remainder of the division of $n$ by $k$.
Combining the result above with Theorem 4.1, we immediately obtain the bidiagonal factorisation of the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials on the step-line as follows.
Corollary 5.2.
@@
The Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials on the step-line, multiple orthogonal with respect to $(\mu_{1},\cdots,\mu_{r})$ given by (5.1), admits the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$, as a product of bidiagonal matrices of the form in (1.2) with nontrivial entries $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ given by (5.8).
To prove the recurrence relation (5.7), we need the following relation for contiguous hypergeometric series, which can be found in [22, Lemma 14.1].
Lemma 5.3.
For $p,q\in\mathbb{N}$, $\vec{c}=\left(c_{1},\cdots,c_{p}\right)\in\mathbb{N}^{p}$, and $\vec{d}=\left(d_{1},\cdots,d_{q}\right)\in\mathbb{N}^{q}$,
${}_{p}F_{q}\left(\left.\begin{matrix}\vec{c}+\vec{e}_{i}\\ \vec{d}+\vec{e}_{j}\end{matrix}\,\right|\,t\right)-{}_{p}F_{q}\left(\left.\begin{matrix}\vec{c}\\ \vec{d}\end{matrix}\,\right|\,t\right)=\,\dfrac{\left(d_{j}-c_{i}\right)\prod\limits_{k=1,\,k\neq i}^{p}c_{k}}{\left(d_{j}+1\right)\prod\limits_{l=1}^{q}d_{l}}\;t\,{}_{p}F_{q}\left(\left.\begin{matrix}\vec{c}+1\\ \vec{d}+1+\vec{e}_{j}\end{matrix}\,\right|\,t\right),$ (5.9)
where $\vec{e}_{i}\in\mathbb{N}^{p}$ and $\vec{e}_{j}\in\mathbb{N}^{q}$ are the vectors whose entries are all 0 except the $i^{\textrm{th}}$-entry, respectively $j^{\textrm{th}}$-entry, which is equal to $1$.
This formula can be easily proved by comparing the coefficients of $t^{n}$ on both sides.
Proof of Theorem 5.1.
For $n=0$, $n_{0}^{\,\prime}=0$ and $n_{i}^{\,\prime}=n_{i}=0$ for all $1\leq i\leq r$. Hence, (5.5) reduces to
$g_{0}(t)=(1-t)^{b}{}_{1}F_{0}\left(\left.\begin{matrix}b\\ -\end{matrix}\,\right|\,t\right)=1.$ (5.10)
For $1\leq n\leq r$, $n_{0}^{\,\prime}=1$, $n_{i}^{\,\prime}=n_{i}=1$ if $1\leq i\leq n-1$, $_{n}^{\,\prime}=0, nn$$_{n}=1$, and $n_{i}^{\,\prime}=n_{i}=0$ if $n+1\leq i\leq r$. As a result, (5.5) reduces to
$g_{n}(t)=(1-t)^{b}{}_{2}F_{1}\left(\left.\begin{matrix}b+1,a_{n}+b\\ a_{n}+b+1\end{matrix}\,\right|\,t\right)={}_{2}F_{1}\left(\left.\begin{matrix}a_{n},1\\ a_{n}+b+1\end{matrix}\,\right|\,t\right)=\sum_{k=0}^{\infty}\frac{\left(a_{n}\right)_{k}}{\left(a_{n}+b+1\right)_{k}}\,t^{k}.$ (5.11)
Here, the second equality is a particular case of [13, Eq. 15.8.1] and the final equality follows from the definition of the hypergeometric series.
Using the vectors $\vec{n}=\left(n_{1},\cdots,n_{r}\right)\in\mathbb{N}^{r}$ and $\vec{n}^{\,\prime}=\left(n_{0}^{\,\prime},\cdots,n_{r}^{\,\prime}\right)\in\mathbb{N}^{r+1}$ defined below (5.5) and introducing the vectors $\vec{a}=\left(a_{1},\cdots,a_{r}\right)\in\mathbb{C}^{r}$ and $\vec{a}^{\,\prime}=\left(0,a_{1},\cdots,a_{r}\right)\in\mathbb{C}^{r+1}$, we can rewrite (5.5) as
$g_{n}(t)=(1-t)^{b}{}_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\vec{n}^{\,\prime}\vskip 2.0pt\\ \vec{a}+b+\vec{n}\end{matrix}\,\right|\,t\right)\quad\text{for all }n\in\mathbb{N}.$ (5.12)
Fix $n\in\mathbb{N}$. To find the recurrence relation (5.7), we start by observing that
$\overrightarrow{n+1}=\vec{n}+\vec{e}_{j(n)}\quad\text{and}\quad\overrightarrow{n+1}^{\,\prime}=\vec{n}^{\,\prime}+\vec{e}_{i(n)}^{\,\prime}$, (5.13)
where $\vec{e}_{j(n)}=\left(e_{j(n),1},\cdots,e_{j(n),r}\right)$ and $\vec{e}_{i(n)}^{\,\prime}=\left(e_{i(n),0}^{\,\prime},\cdots,e_{i(n),r}^{\,\prime}\right)$ are vectors whose entries are all equal to 0, except $e_{i(n),i(n)}^{\,\prime}=1=e_{j(n),j(n)}$. Therefore, we have
$g_{n+1}(t)-g_{n}(t)=(1-t)^{b}\left({}_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\vec{n}^{\,\prime}+\vec{e}_{i(n)}^{\,\prime}\vskip 2.0pt\\ \vec{a}+b+\vec{n}+\vec{e}_{j(n)}\end{matrix}\,\right|\,t\right)-{}_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\vec{n}^{\,\prime}\vskip 2.0pt\\ \vec{a}+b+\vec{n}\end{matrix}\,\right|\,t\right)\right).$ (5.14)
Using Lemma 5.3,
$_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\vec{n}^{\,\prime}+\vec{e}_{i(n)}^{\,\prime}\vskip 2.0pt\\ \vec{a}+b+\vec{n}+\vec{e}_{j(n)}\end{matrix}\,\right|\,t\right)-{}_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\vec{n}^{\,\prime}\vskip 2.0pt\\ \vec{a}+b+\vec{n}\end{matrix}\,\right|\,t\right)=\alpha_{n}\,t\,{}_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\vec{n}^{\,\prime}+1\vskip 2.0pt\\ \vec{a}+b+\vec{n}+\vec{e}_{j(n)}+1\end{matrix}\,\right|\,t\right),$ (5.15)
with
$\alpha_{n}=\dfrac{\left(a_{j(n)}-a_{i(n)}+n_{j(n)}-n_{i(n)}^{\,\prime}\right)\prod\limits_{i=0,\,i\neq i(n)}^{r}\left(a_{i}+b+n_{i}^{\,\prime}\right)}{\left(a_{j(n)}+b+n_{j(n)}+1\right)\prod\limits_{j=1}^{r}\left(a_{j}+b+n_{j}\right)}.$ (5.16)
Observe that $\vec{n}^{\,\prime}+1=\overrightarrow{n+r+1}^{\,\prime}$ and $\vec{n}+\vec{e}_{j(n)}+1=\overrightarrow{n+1}+1=\overrightarrow{n+r+1}$. Hence,
$g_{n+1}(t)-g_{n}(t)=\alpha_{n}\,t\,(1-t)^{b}\,{}_{r+1}F_{r}\left(\left.\begin{matrix}\vec{a}^{\,\prime}+b+\overrightarrow{n+r+1}^{\,\prime}\vskip 2.0pt\\ \vec{a}+b+\overrightarrow{n+r+1}\end{matrix}\,\right|\,t\right)=\alpha_{n}\,t\,g_{n+r+1}(t).$ (5.17)
Finally, note that
$n_{j}=\begin{cases}\left\lfloor\frac{{n}}{r}\right\rfloor+1&\text{if }j\leq j(n)-1,\vskip 2.0pt\\ \hfil\left\lfloor\frac{{n}}{r}\right\rfloor&\text{if }j\geq j(n),\end{cases}\quad\text{and}\quad n_{i}^{\,\prime}=\begin{cases}\left\lfloor\frac{{n}}{r+1}\right\rfloor+1&\text{if }i\leq i(n)-1,\vskip 2.0pt\\ \hfil\left\lfloor\frac{{n}}{r+1}\right\rfloor&\text{if }i\geq i(n).\end{cases}$ (5.18)
Therefore, (5.16) is equivalent to (5.8), concluding our proof. ∎
Here, we find again the bidiagonal factorisation described in Corollary 5.2 for the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials, now using explicit expressions for the type I polynomials recently given in [6]. We consider the system of measures $\left(\mu_{1},\cdots,\mu_{r}\right)$ supported on the interval $(0,1)$ with densities $\mathrm{d}\mu_{j}(x)=x^{a_{j}-1}(1-x)^{b}\,\mathrm{d}x$, with b>1b>-1 and $a_{1},\cdots,a_{r}>0$ such that $i\neq j$ implies $a_{i}-a_{j}\not\in\mathbb{Z}$. Then, for any $1\leq j\leq r$, we have:
  • $\mu_{j}\left(a_{1},\cdots,a_{r};b\right)=\mu_{1}\left(a_{j},\cdots,a_{r},a_{1}+1,\cdots,a_{j-1}+1;b\right)$,
  • $x\,\mu_{j}\left(a_{1},\cdots,a_{r};b\right)=\mu_{j}\left(a_{1}+1,\cdots,a_{r}+1;b\right)=\mu_{1}\left(a_{j}+1,\cdots,a_{r}+1,a_{1}+2,\cdots,a_{j-1}+2;b\right)$.
Therefore, the multiple orthogonal polynomials with respect to the system $\left(\mu_{i+1},\cdots,\mu_{r},x\,\mu_{1},\cdots,x\,\mu_{i}\right)$, $0\leq i\leq r$, are the Jacobi-Piñeiro polynomials with parameters $a_{i},\cdots,a_{r},a_{1}+1,\cdots,a_{i-1}+1$, and $b$.
As a result, it follows from [6, Thm. 2] that, in this case, the coefficients of the type I multiple orthogonal polynomials appearing in (3.18)-(3.19) are, for $m\in\mathbb{N}$ and $0\leq k\leq r-1$,
$a^{[i]}_{rm+k,k+1}(m)=\frac{\Gamma\left(a_{k+i}+b+(r+1)m+k\right)\prod\limits_{j=1}^{k}\left(a_{j+i}+b+rm+k\right)_{m+1}\prod\limits_{j=k+1}^{r}\left(a_{j+i}+b+rm+k\right)_{m}}{m!\,\Gamma\left(a_{k+i}+m\right)\Gamma(b+rm+k)\prod\limits_{j=1}^{k-1}\left(a_{k+i+1}-a_{j+i}\right)_{m+1}\prod\limits_{q=k+1}^{r}\left(a_{k+i+1}-a_{j+i}+1\right)_{m}},$ (5.19)
where $a_{rn+j}=a_{j}+n$ for all $0\leq j\leq r$ and $n\in\mathbb{N}$.
Inputting these coefficients in (3.18)-(3.19), we derive the formulas
$\alpha_{(r+1)(rm+k)}=\frac{(a_{k+1}+m)\prod\limits_{j=1}^{r}\left(a_{j}+b+rm+k\right)}{\prod\limits_{j=1}^{r+1}\left(a_{k+j}+b+(r+1)m+k\right)},$ (5.20)